### v2

Abstract
Large Multimodal Models (LMMs) demonstrate impressive capabilities. However, current benchmarks predominantly focus on image comprehension in specific domains, and these benchmarks are labor-intensive to construct. Moreover, their answers tend to be brief, making it difficult to assess the ability of LMMs to generate detailed descriptions of images. To address these limitations, we propose the MMGenBench-Pipeline, a straightforward and fully automated evaluation pipeline. This involves generating textual descriptions from input images, using these descriptions to create auxiliary images via text-to-image generative models, and then comparing the original and generated images. Furthermore, to ensure the effectiveness of MMGenBench-Pipeline, we design MMGenBench-Test, evaluating LMMs across 13 distinct image patterns, and MMGenBench-Domain, focusing on generative image performance. A thorough evaluation involving over 50 popular LMMs demonstrates the effectiveness and reliability of both the pipeline and benchmark. Our observations indicate that numerous LMMs excelling in existing benchmarks fail to adequately complete the basic tasks related to image understanding and description. This finding highlights the substantial potential for performance improvement in current LMMs and suggests avenues for future model optimization. Concurrently, MMGenBench-Pipeline can efficiently assess the performance of LMMs across diverse domains using only image inputs.
{strip}[Uncaptioned image]
Figure 1: The MMGenBench-Test consists of $13$ distinct image patterns, each of which includes several images. The text, accompanied by a corresponding pattern, serves as a concise explanation of that specific image pattern. Please refer to the Appendix B.1 for more details.
1 Introduction
We have witnessed rapid progress of Large Multimodal Models (LMMs) [42, 81, 27, 66, 50, 4, 6, 12, 65, 10, 53], which efficiently utilize the strength of LLMs [54, 61, 53, 68, 32, 52] in processing visual and textual inputs. Compared to text, visual images are characterized by their high level of abstraction and information density, while exhibiting strong spatial correlations and structural complexity. The development of comprehensive benchmarks is crucial for enhancing and accurately evaluating LMMs. Specifically, many popular benchmarks [41, 30, 76, 72, 80, 17, 19, 45, 55], provide standardized evaluations for LMMs by assessing multimodal tasks across various datasets. However, these benchmarks frequently depend on traditional datasets, resulting in issues of data leakage and limited task diversity. Although many various tasks such as VQA [2, 49, 79, 51, 47, 24, 55], Image Caption [33, 44], and OCR [77, 45, 18, 74] are carefully designed, as shown in Fig. 2(a), most existing benchmarks require expensive manual construction and focus on specific domains, making it difficult to extend to other domains. The constructed sample answers are mostly brief, focusing only on evaluating the understanding performance of LMMs and ignoring the evaluation of the ability to generate detailed descriptions of images. In addition, the understanding and generation of images remain disparate fields, with the most powerful models [16, 34, 53] in their respective domains adhering to distinct paradigms. For instance, GPT-4 [53], which is grounded in the next token prediction paradigm, exhibits an impressive capacity for image comprehension. Similarly, Flux [34] has achieved noteworthy success in text-to-image synthesis by leveraging diffusion models [25, 58]. This divergence highlights the complexity of achieving a unified approach to image processing and synthesis, as the state-of-the-art techniques continue to evolve along separate trajectories. Furthermore, LMMs are extensively employed to generate data for generative models [56, 70, 39, 28]. It is noteworthy that LMMs excel in image-to-text tasks, while diffusion models are particularly effective in text-to-image tasks. A robust understanding of an image implies that LMMs can distill its essential information into text prompts, which text-to-image models can use to reconstruct the scene to a certain extent. This process can be viewed as a form of “compression”. Hence, it is both reasonable and significant to evaluate the performance of LMMs using diffusion models.
In this paper, we propose MMGenBench-Pipeline, a fully automated evaluation pipeline (refer to Fig. 2(b)) that initially allows LMMs to generate textual descriptions from input images, then employs text-to-image generative models to create auxiliary images. Finally, we use an image representation model to obtain the embeddings of images and perform post-processing to assess the performance of LMMs in image understanding and description. To verify the effectiveness of MMGenBench-Pipeline, we introduce MMGenBench-Test, a comprehensive benchmark designed to evaluate LMMs across 13 distinct image patterns (refer to Fig. 1), and MMGenBench-Domain, which focuses on assessing LMMs performance within the generative image domain. Extensive experiments on over 50 popular LMMs demonstrate the effectiveness and reliability of both the pipeline and benchmark. Notably, our findings reveal that numerous LMMs excelling in existing benchmarks fail to address the basic tasks of image understanding and description, highlighting the substantial potential for improvement and optimization in future models. The MMGenBench-Pipeline also enables efficient evaluation of LMMs across diverse domains through image inputs alone, providing a flexible and scalable benchmarking tool.
In summary, our contributions are as follows:
  • Fully Automated Evaluation Pipeline: We propose the first fully automated pipeline MMGenBench-Pipeline designed to evaluate the capabilities of LMMs in image understanding and description by solely utilizing images. This pipeline utilizes text-to-image models and image representation models for automated evaluation, thereby markedly minimizing human involvement and improving the efficiency and objectivity of the evaluation procedure.
  • Comprehensive Benchmarks: In order to verify the effectiveness of MMGenBench-Pipeline, we developed MMGenBench-Test, a comprehensive benchmark designed to evaluate LMMs across $13$ image patterns, and MMGenBench-Domain, which assesses the performance of LMMs in the generative image domain.
  • Extensive Evaluation: Our study includes a broad evaluation of over 50 popular LMMs, providing critical insights into their capabilities and limitations in basic image understanding and description tasks.
Please refer to Appendix A for related work.
Figure 2: Comparison between previous benchmarks and MMGenBench. MMGenBench has several novel features: 1) Based on powerful text-to-image models and image representation models, MMGenBench can fully automatically complete the evaluation of LMMs without the need for expensive manual annotation; 2) MMGenBench can easily evaluate the performance of LMMs in any domain, whereas previous benchmarks could mostly only evaluate the performance in specific domains; 3) The “answer” to previous benchmarks were mostly brief, overlooking the basic ability to generate detailed descriptions of images.
Figure 3: An overview of the MMGenBench-pipeline, illustrating the fully automated evaluation process. It starts by receiving user input (including the task instruction prompt and input images), and then generates the corresponding textual descriptions of input images. Subsequently, this process is followed by using a powerful text-to-image model to generate auxiliary images, then produces the representation of the input images and the generated ones using an image representation model, and finally outputs the evaluation score of LMMs.
2 The MMGenBench-Pipeline
2.1 Fully Automated Evaluation Pipeline
The proposed pipeline for LMMs, based on text-to-image generative models and image representation models, consists of four components: test set construction, textual description generation, auxiliary image generation, and quantitative metric computation.
Test Set Construction. Our method can easily evaluate image understanding and description capabilities of LMMs in any domain. As shown in Stage 1 of Fig. 3, given any image from the domain to be tested, we can construct the multi-modal input for the LMMs to be evaluated simply using the predefined prompts, and apply it to the subsequent process.
Textual Description Generation. The input modalities for the LMM primarily consist of two parts: the image used for model understanding and inference, and the prompt guiding the inference direction. To facilitate a standardized workflow, we employ a manually crafted, normalized prompt $\mathbf{P}_{art}$. This prompt constrains the LMM’s comprehension and reasoning across five dimensions: role, definition explanation, task instruction, key points and requirements, and output format (see Appendix C.1 for details). This process can be formalized as follows:
$\mathbf{P}_{gen}=\text{LMM}(\mathbf{I}_{inp},\mathbf{P}_{art}).$ (1)
Through the above process, we obtain the detailed textual description of an image $\mathbf{P}_{gen}$, generated by the LMM, which will serve as the input for the subsequent stage.
Auxiliary Image Generation. In this stage, the main process involves the utilization of text-to-image models to generate auxiliary images based on the given textual descriptions. Theoretically, the generation quality is highly dependent on the chosen text-to-image model. To minimize variable interference, we standardize the evaluation by selecting four state-of-the-art models: FLUX.1-dev [34], Stable Diffusion 3.5 [16], Kolors [62] and Lumina [20]. The comparison results across these models provide cross-validation of generation effectiveness. This phase is represented as:
$\mathbf{I}_{gen}=G(\epsilon;\mathbf{P}_{gen},\theta),$ (2)
where $\mathbf{I}_{gen}$ denotes the generated image, $\epsilon$ represents a randomly sampled variable from the latent space, typically following a certain distribution (e.g., Gaussian distribution), $\theta$ represents the model parameters, and $G$ denotes the image generation function.
Quantitative Metric Computation. We quantify the functionality of LMM by evaluating the similarity between the input image $\mathbf{I}_{inp}$ and the generated one $\mathbf{I}_{gen}$. Since most generative models introduce a certain degree of randomness in their inference process, achieving pixel-level consistency between images is challenging. To address this, we conduct comparisons at the representational level. Specifically, we utilize the Unicom [5] model to encode each image and obtain its representations. This phase is encapsulated as:
$\mathbf{F}_{I_{inp}}=\text{Encoder}(\mathbf{I}_{inp}),\mathbf{F}_{I_{gen}}=\text{Encoder}(\mathbf{I}_{gen}),$ (3)
where $\mathbf{F}_{I_{inp}}$ and $\mathbf{F}_{I_{gen}}$ represent the features extracted from the input image $\mathbf{I}_{inp}$ and the generated image $\mathbf{I}_{gen}$, respectively. To provide a quantitative evaluation, we compute both a similarity score and a generation quality score based on these feature representations.
2.2 Evaluation Metric
SIM-Score is a metric for evaluating the similarity between two features, commonly using cosine similarity. In this study, SIM-Score is calculated as follows:
[…]
The FID Score (FID-Score) measures the difference between the distributions of generated images and real images. A lower FID Score indicates that the generated images have higher quality and greater similarity to the real images. The calculation is primarily based on the below equation:
$\text{FID}=\|\mu_{x}-\mu_{y}\|^{2}+\text{Tr}(\Sigma_{x}+\Sigma_{y}-2(\Sigma_{x}\Sigma_{y})^{1/2}),$ (5)
where $\mu_{x}$ and $\mu_{y}$ denote the means of the input image features $F_{I_{inp}}$ and the generated image features $F_{I_{gen}}$, respectively, while $\Sigma_{x}$ and $\Sigma_{y}$ represent their covariances. The notation $\text{Tr}(\cdot)$ denotes the trace of a matrix.
3 The MMGenBench Benchmark
3.1 Overview
Although the existing benchmarks can evaluate the image understanding capabilities of LMMs from various dimensions, as mentioned in Sec. 1, they still lack the evaluation of the basic image understanding of LMMs and the ability to describe the images clearly.
To effectively measure the understanding and description capabilities of LMMs across various types of images, we constructed a high-quality test set MMGenBench-Test for $13$ image patterns using the JourneyDB [60] test set. We proposed a multi-stage method for extracting and annotating image patterns as illustrated in Fig. 5. To ensure the results’ accuracy, we manually double-checked the image patterns and performed the final annotations.
In addition, we have also constructed a dataset in the “image generation” domain, termed MMGenBench-Domain, to evaluate the ability of LMMs in the understanding and describing “generated images” task. It is important to emphasize that MMGenBench-Pipeline can measure the ability of LMMs to understand and generate detailed image descriptions in any domain. By utilizing images from a particular domain, we can easily assess the performance of LMMs specific to that domain.
Figure 4: Statistics of MMGenBench-Test, which contains $13$ image patterns with $1,284$ images. More details are in Sec. 3.
3.2 Dataset Statistics
In the MMGenBench-Test dataset, we constructed a high-quality test set containing $1,284$ images across $13$ distinct image patterns (see Fig. 1). The distribution of images per pattern is shown in Fig. 4, which illustrates that each pattern contains a minimum of $114$ images. Please note that an image may contain multiple patterns. For instance, the first image annotation in Fig. 5 contains four image patterns: “Surreal”, “Natural”, “Artistic” and “Color”. These $13$ image patterns are carefully designed so that the dataset can measure the image comprehension and description capabilities of LMMs across diverse dimensions. Additionally, all samples in the dataset are manually checked and annotated to ensure their overall quality.
To construct MMGenBench-Domain, we randomly sampled $10,000$ images from the JourneyDB validation set. By utilizing the MMGenBench-pipeline, we can evaluate the image understanding and descriptive performance of LMMs within this domain, obviating the need for additional data.
Figure 5: An overview of the MMGenBench-Test benchmark construction process. We first use GPT-4o to extract the image patterns from the input images. Then, we use GPT-4 Turbo to summarize these patterns and manually select $13$ patterns. Subsequently, GPT-4o is employed again to re-annotate these patterns. These annotations are reviewed and modified to produce the final result by human annotators.
3.3 Data Collection
To construct MMGenBench-Test, we first extract all images from the JourneyDB test set and process them using the process shown in Fig. 5. For the domain-specific dataset, we use the JourneyDB validation set for its construction. Subsequently, we will elaborate on each processing step.
Pattern Extraction. We extract image patterns from existing images to measure the understanding and description ability of LMMs across various image categories. Given that we possess only images but no additional information, we utilize the powerful GPT-4o model to extract and analyze the underlying image patterns. The task is executed by GPT-4o in three sequential steps: 1) providing a detailed description of the image; 2) annotating the possible patterns based on the image features and description; and 3) explaining the rationale behind the annotated image patterns. By utilizing our carefully designed prompts (refer to Appendix B.3), GPT-4o can perform the task effectively. An example is presented in Fig. 10 in Appendix B.2, wherein the image description, alongside the extracted image patterns and rationales, aligns accurately with the image content. By annotating each image in the JourneyDB test set, we ultimately obtained a total of $1868$ image patterns.
Annotation Summary. By leveraging GPT-4o, we extracted an extensive range of image patterns and observed that the semantic meanings across different patterns may be consistent. We conducted thorough statistics of the image patterns, by quantifying the occurrences of each type and ranking them in descending order (as seen in Appendix B.1). Subsequently, GPT-4-Turbo was utilized to generate the summary. To ensure the accuracy and validity of the extracted patterns, we carefully designed model prompts for this task (shown in Appendix B.3). This procedure ultimately resulted in a hand-crafted list of $13$ distinct image patterns, each thoroughly described in Appendix B.1.
Image Re-annotation. We conducted a re-annotation of the JourneyDB test set using the previously identified $13$ image patterns. Similar to Image Pattern Extraction, GPT-4o was employed to generate descriptions for each image. Subsequently, these descriptions were annotated based on the predefined image patterns, accompanied by the underlying reasons. A key difference in this process is that re-annotation is restricted to the specified $13$ image patterns, excluding any other patterns. Given the extensive number of images in the test set, it is essential to select a subset of the images for subsequent operations. The re-annotated images were systematically classified into distinct image patterns, each containing a specified number of images. To construct the final dataset, we randomly sampled a subset of images from each pattern with a selection probability of $\frac{100}{N}$, where $N$ represents the total number of images within the pattern. To mitigate potential annotation errors, we manually verified and annotated each image. The final dataset was established by voting between the model predictions and annotations.
Domain Data Collection. Constructing high-quality datasets is an extremely expensive task. Therefore, it is crucial to develop an efficient and convenient method for creating domain-specific datasets when evaluating the image understanding and description capabilities of LMMs in a new domain. By leveraging the MMGenBench-pipeline, our method enables a seamless evaluation of LMM performance across various domains, only depending on the availability of domain-specific images. In this study, we randomly select $10,000$ images from the JourneyDB validation set to create our domain-specific dataset. To quantify LMM performance within this domain, we calculate both FID-Score and SIM-Score.
4 Experiment
Figure 6: The comparative analysis of four different text-to-image models on MMGenBench. The horizontal axis denotes the index of LMMs. Fig. (a) illustrates the SIM-Score for over 50 LMMs on MMGenBench-Test, while Fig. (b) presents the FID-Score for the same set of models on MMGenBench-Test.
4.1 Experimental Setup
Evaluation Models. In the evaluation of LMMs, we selected over 50 models that exhibited strong performance on the VLM leaderboard [15]. We mainly focus on open-source models, which include Qwen2-VL [65], InternVL2 [10], LLaVA-OV [37], Ovis [50], etc. Additionally, we evaluated three closed-source API models: GPT-4o [53], Qwen-VL-Max [6], and Qwen-VL-Plus [6]. To process the textual descriptions generated by LMMs and create auxiliary images, we selected four text-to-image generation models: FLUX.1-dev [34], Stable Diffusion 3.5 [16], Kolors [62] and Lumina [20]. During the final evaluation phase, the Unicom [5] image representation model was utilized to extract image features. Further details about our baseline models are provided in the Appendix C.2.
Implementation Details. During the textual descriptions generation process in LMMs, we utilize the default settings of VLMEvalKit [15], modifying only the predefined query prompt and the image to meet the task-specific requirements. Subsequently, we employ four distinct text-to-image models to generate auxiliary images and extract features using an image representation model. We then compute both the SIM-Score and FID-Score for evaluation. We analyzed the performance metrics of four text-to-image models on the MMGenBench-Test dataset, as illustrated in Fig. 6. The results indicate that the four text-to-image models exhibited consistency in both SIM-Score and FID-Score. Therefore, unless otherwise stated, the reported results are obtained using the FLUX.1-dev text-to-image model. Please refer to Appendix C.2 for more detailed results.
Model Test Domain
[…]
GPT-4o [53] 0.566 1.306 - -
Qwen-VL-Max [6] 0.552 1.363 - -
Qwen-VL-Plus [6] 0.475 1.586 - -
Qwen2-VL-72B [65] 0.553 1.357 0.545 0.710
Qwen2-VL-7B [65] 0.532 1.437 0.524 0.775
Qwen2-VL-2B [65] 0.487 1.549 0.501 0.806
InternVL2-76B [10] 0.599 1.264 0.599 0.632
InternVL2-40B [10] 0.566 1.350 0.566 0.696
InternVL2-26B [10] 0.576 1.320 0.577 0.671
InternVL2-8B [10] 0.547 1.403 0.548 0.701
InternVL2-4B [10] 0.556 1.345 0.556 0.689
InternVL2-2B [10] 0.476 1.563 0.483 0.848
Ovis1.6-Gemma2-9B [50] 0.582 1.316 0.579 0.667
Ovis1.5-Gemma2-9B [50] 0.521 1.487 0.524 0.808
Ovis1.5-Llama3-8B [50] 0.526 1.466 0.527 0.795
LLaVA-OV-72B [37] 0.494 1.561 0.491 0.872
LLaVA-OV-SI-72B [37] 0.512 1.512 0.514 0.813
LLaVA-OV-7B [37] 0.488 1.590 0.490 0.861
LLaVA-OV-SI-7B [37] 0.492 1.554 0.497 0.834
MiniCPM-V2.6-8B [75] 0.548 1.386 0.545 0.710
RBDash-72B [57] 0.525 1.413 0.527 0.740
xGen-MM-4.4B [73] 0.414 1.671 0.412 0.940
MiniCPM-V2.5-8B [75] 0.526 1.432 0.530 0.767
Table 1: Experiment results of different LMMs on MMGenBench-Test/Domain, where SIM corresponds to SIM-Score, FID corresponds to FID-Score. More details are in Sec. C.3.
4.2 Main Results
Overall Performance. As shown in Table 1 and 2, the SIM-Score of the most advanced LMMs on MMGenBench-Test is below $0.600$. Specifically, GPT-4o achieved a score of $0.566$, which is lower than the $0.599$ obtained by the open-source model InternVL2-76B. Notably, there is no clear correlation between the model size and performance across different series of models. This indicates the importance of training data and the training process in developing LMMs with strong capabilities of image understanding and description. Furthermore, it is observed that models excelling on existing benchmarks do not necessarily perform well on MMGenBench-Test, such as LLaVA-OV, which only scores $0.494$. In the MMGenBench-Domain dataset, the SIM-Score results align closely with those of the MMGenBench-Test. However, there is a significant difference in FID-Score because MMGenBench-Domain includes 10,000 images, thereby improving the accuracy of its FID-Score measurement. Therefore, we propose using SIM-Score as the primary metric. Detailed results and comprehensive analysis are provided in the following sections as well as in Appendix C.3.
Model [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
GPT-4o [53] 0.581 0.543 0.516 0.546 0.576 0.574 0.533 0.581 0.567 0.558 0.586 0.620 0.568
Qwen-VL-Max [6] 0.568 0.520 0.506 0.536 0.573 0.558 0.521 0.554 0.553 0.541 0.572 0.608 0.538
Qwen-VL-Plus [6] 0.497 0.455 0.440 0.386 0.516 0.486 0.457 0.444 0.467 0.464 0.496 0.484 0.486
Qwen2-VL-72B [65] 0.572 0.528 0.502 0.525 0.578 0.559 0.526 0.562 0.563 0.549 0.567 0.603 0.548
Qwen2-VL-7B [65] 0.550 0.492 0.480 0.514 0.556 0.539 0.488 0.540 0.535 0.515 0.542 0.584 0.520
Qwen2-VL-2B [65] 0.505 0.462 0.439 0.456 0.519 0.489 0.456 0.486 0.495 0.475 0.494 0.539 0.486
InternVL2-76B [10] 0.616 0.572 0.540 0.597 0.614 0.611 0.569 0.604 0.594 0.581 0.608 0.640 0.588
InternVL2-40B [10] 0.584 0.527 0.518 0.528 0.592 0.577 0.528 0.581 0.563 0.560 0.574 0.607 0.565
InternVL2-26B [10] 0.596 0.547 0.522 0.562 0.586 0.583 0.538 0.587 0.572 0.569 0.581 0.627 0.564
InternVL2-8B [10] 0.564 0.522 0.501 0.526 0.576 0.557 0.507 0.539 0.545 0.536 0.560 0.596 0.543
InternVL2-4B [10] 0.573 0.524 0.507 0.534 0.578 0.564 0.525 0.540 0.546 0.540 0.574 0.587 0.552
InternVL2-2B [10] 0.501 0.450 0.440 0.395 0.511 0.479 0.440 0.469 0.479 0.453 0.501 0.498 0.471
Ovis1.6-Gemma2-9B [50] 0.601 0.547 0.518 0.573 0.593 0.593 0.540 0.595 0.579 0.568 0.580 0.633 0.576
Ovis1.5-Gemma2-9B [50] 0.541 0.493 0.470 0.490 0.546 0.526 0.483 0.541 0.535 0.515 0.532 0.557 0.528
Ovis1.5-Llama3-8B [50] 0.540 0.505 0.485 0.499 0.543 0.526 0.489 0.530 0.539 0.518 0.540 0.565 0.520
LLaVA-OV-72B [37] 0.512 0.463 0.441 0.463 0.518 0.500 0.458 0.504 0.494 0.478 0.510 0.544 0.509
LLaVA-OV-SI-72B [37] 0.530 0.477 0.457 0.500 0.540 0.518 0.467 0.512 0.517 0.515 0.527 0.559 0.509
LLaVA-OV-7B [37] 0.504 0.452 0.439 0.470 0.512 0.496 0.432 0.491 0.491 0.465 0.513 0.541 0.478
LLaVA-OV-SI-7B [37] 0.514 0.462 0.439 0.454 0.520 0.498 0.446 0.494 0.504 0.485 0.514 0.543 0.485
MiniCPM-V2.6-8B [75] 0.559 0.517 0.492 0.524 0.565 0.559 0.500 0.577 0.540 0.547 0.558 0.594 0.522
RBDash-72B [57] 0.533 0.499 0.491 0.497 0.545 0.531 0.506 0.549 0.540 0.514 0.533 0.553 0.512
xGen-MM-4.4B [73] 0.420 0.399 0.391 0.289 0.464 0.414 0.402 0.409 0.412 0.430 0.468 0.438 0.415
MiniCPM-V2.5-8B [75] 0.539 0.498 0.478 0.485 0.547 0.534 0.496 0.523 0.529 0.526 0.539 0.555 0.521
Table 2: Experimental results on MMGenBench-Test, which demonstrate the SIM-Score on each image pattern. Information about the icons in the first row and more detailed results can be found in Fig. 1, Appendix B.1 and C.3.
Figure 7: Model performance by image patterns. Please refer to Table 2 and Sec. 4.2 for more results and discussions.
Figure 8: The comparison of SIM-Score with different LMMs on MMGenBench-Test: (a) LMMs from the same series but with different parameter sizes; (b) LMMs with the same sizes but using different training protocol or data. More details are in Sec. 4.2.
Image Patterns Revealing Model’s Strengths and Limitations. Fig. 7 shows the accuracy of the best-performing LMMs on the MMGenBench dataset. We observed that LMMs exhibit superior performance on image patterns such as “Artistic”, “Surreal”, “Symbol”, “Color”, etc. Conversely, their performance declines on patterns such as “Contextual”, “Orientation”, “Count”, “Motion”, etc. This discrepancy suggests that LMMs are proficient in tasks requiring coarse-grained image understanding and description. In contrast, their abilities for fine-grained understanding and description are more challenging, as these tasks require the understanding of the complex contextual relationships within images and the precise linguistic description.
Impact of Parameter Scales, Training Methods, and Data Variations. Fig.8(a) shows the performance results of the same series of models with varying parameter sizes, while Fig.8(b) presents the performance results of the same model and parameters under different training protocols and datasets. We can see that as the model parameters increase, the performance of Qwen2-VL improves from $0.487$ to $0.553$, and the results of InternVL2 increased from $0.476$ to $0.599$. Additionally, by using improved training data and protocols, Ovis1.6 outperforms Ovis1.5 by $0.06$. Given that our evaluation is based on a single input image, LLaVA-OV-SI demonstrates superior performance to LLaVA-OV.
Figure 9: Qualitative Results on MMGenBench. We present common problems identified in the experiments, which include issues with output format, generated content, and final results. Please refer to Sec. 4.3 and Appendix D for more discussions and results.
4.3 Analysis
We identify three common problems that significantly affect the performance of LMMs in image understanding and description. Please refer to the Appendix D for more details.
Failure in Following Instructions. As shown in Fig. 9, we have selected various LMMs to demonstrate a portion of the qualitative examples. It is observed that certain LMMs fail to strictly follow instructions when generating the textual descriptions. Specifically, some LMMs prepend a prefix to the textual descriptions (e.g., RBDash-72B [57] in Fig. 9), while others append explanatory content following the textual descriptions (e.g., MiniCPM-Llama3-V2.5 [75] in Fig. 9). Additionally, it is also evident that the ability to follow instructions does not appear to be correlated with the parameter size of different series LMMs. For example, both Qwen2-VL-2B and InternVL2-2B accurately generate the textual descriptions following the given instructions. We argue that an effective LMM, especially when undergoing instruction tuning, should be capable of accurately following instructions and successfully completing the task.
Inability in Generating Detailed Textual Descriptions. Most of the image-text pairs used in the training of existing LMMs have relatively short image captions. Consequently, numerous models exhibit suboptimal performance on the MMGenBench Benchmark, which requires the model to understand the image and provide a detailed description. As shown in Fig. 9, LLaVA-OV-72B [37], despite its extensive 72B parameters, produces relatively short textual descriptions, which hinders its ability to excel on the MMGenBench Benchmark. Additionally, we observed that for LMMs in the same series, smaller LMMs tend to generate shorter textual descriptions, as evidenced in the Qwen2-VL and InternVL2 series. Based on these findings, we hypothesize that incorporating image-description pairs during the training phase of LMMs could achieve superior results.
Model Overfitting. Existing LMMs frequently undergo task-specific training to meet established benchmarks. This specialized training can lead to overfitting, which may compromise the basic image understanding and description capabilities of LMMs. For instance, the xGen-MM [73] results in Fig. 9 demonstrate an overfitting towards the notion of “safety”, despite the given image not presenting any safety risk. This observation suggests that in adapting models to various tasks, it is crucial to maintain the basic image understanding and description capabilities of LMMs.
5 Conclusion
In this paper, we present a straightforward and fully automated evaluation pipeline to comprehensively evaluate the ability of LMMs in image understanding and description. Based on MMGenBench-pipeline, we introduce MMGenBench-Test and MMGenBench-Domain to evaluate LMM performance across different image patterns and domain-specific images. By evaluating over 50 widely used LMMs, we demonstrate the reliability and effectiveness of both the pipeline and benchmarks. In conclusion, our method provides a more fundamental evaluation of existing LMMs, identifies specific gaps in their image understanding and description capabilities, and establishes a robust foundation for further research and model improvement in these areas. All code and data will be released soon.
Supplementary Material
Appendix A Related Work
A.1 Automatic Benchmarks
The rapid advancements in LLMs have prompted the development of diverse benchmarks to evaluate their performance. Early efforts, such as LMExamQA  [7], introduced the ”Language-Model-as-an-Examiner” framework to provide a scalable and comprehensive evaluation for LLMs. However, there remains a significant gap in the evaluation of the reasoning capabilities of LLMs, especially in dynamic contexts. To address this deficiency, recent frameworks like DYVAL  [83] and DYVAL2  [84] have focused on dynamic reasoning tasks. DYVAL concentrates on reasoning abilities, while DYVAL2 extends the evaluation to a broader psychometric approach, thereby providing deeper insights into cognitive evaluations. Additionally, AutoBencher  [38] has automated the generation of challenging and novel datasets, specifically designed for the evaluation of LLM. Platforms such as UNIGEN  [69] and Task Me Anything  [80] aim to further enhance evaluation by developing domain-specific benchmarks tailored to the unique capabilities of LLMs.
A.2 LMM Benchmarks
The emergence of LMMs has highlighted the insufficiency of traditional benchmarks, which primarily focus on isolated task performance and fail to capture the full capabilities exhibited by multimodal models. Early multimodal benchmarks, such as those introduced by [41, 23, 79], established essential foundations but lacked the granularity required for robust evaluation of the model’s understanding and reasoning abilities. Recent studies [46, 9, 78] emphasize the critical need for comprehensive, fine-grained benchmarks that can more accurately evaluate the complex capabilities of LMMs. However, existing benchmarks, such as LVLM-eHub [72] and LAMM [76], still rely on classical datasets that do not fully reflect the current state of the field, and neglect issues such as data leakage during model training. Addressing these concerns, MMStar [9] introduces an innovative approach that eliminates visual content dependencies and mitigates data leakage risks, thereby providing a more reliable and secure evaluation framework. Furthermore, the development of automated benchmarking systems, such as AUTOBENCH-V[8], enhances the scalability and objectivity of the evaluations, by automating the benchmarking process and minimizing human bias. Despite these significant advancements, further refinement is needed to create truly comprehensive and dynamic benchmarks that can capture the evolving capabilities of LMMs.
Appendix B Details of MMGenBench Benchmark
B.1 Details of MMGenBench-Test Data
Thirteen Image Patterns. As illustrated in Fig. 11, summarization is performed using GPT-4 Turbo, followed by human verification. We identify $13$ distinct image patterns, each with a comprehensive explanation. It should be noted that a single image can exhibit multiple patterns concurrently.
Patterns Extracted by GPT-4o. Fig. 12 shows the image patterns extracted by GPT-4o, organized in descending order. While the total number of extracted patterns is 1868, only the most frequently occurring patterns are presented due to space constraints.
B.2 Case Study of Pattern Extraction
GPT-4o has demonstrated strong capabilities in extracting image patterns from given images. Fig. 10 illustrates an instance where GPT-4o effectively identifies and extracts patterns from an image. The results demonstrate a high degree of correspondence between the image description, the extracted pattern, and the corresponding explanation, all of which align precisely with the visual content.
B.3 Details of Construction Prompts
To construct the MMGenBench-Test, a series of prompts have been tailored to the specific requirements of our task meticulously. Fig. 13, 14, and 15 include the three sequential stages in the pipeline (see Fig. 3) used to develop MMGenBench-Test.
Appendix C Additional Experimental Details
C.1 Evaluation Pipeline Prompt
We meticulously crafted a prompt to guide LMMs in generating textual descriptions of images for subsequent evaluation. As shown in Fig. 16, the prompt encompasses five key perspectives: “role”, “definition”, “task”, “key points”, and “output”. To ensure the objectivity and consistency of the evaluation, we employed this singular prompt, as it is sufficiently clear to enable the model to effectively perform the textual description generation task.
C.2 Experimental Setup
We employed over 50 popular LMMs, as detailed in Tables 3 and 4, which are available on the OpenVLM Leaderboard [13]. For open-source LMMs, we utilized the original configuration of VLMEvalKit [15] to perform inference of textual descriptions of images on $8$ NVIDIA-H20 GPUs. During the image generation and metric calculation stages, we still use $8$ NVIDIA-H20 GPUs for multi-process inference to efficiently complete the task.
C.3 Comprehensive Experimental Evaluation
A comprehensive set of experimental results encompassing over 50 popular LMMs is shown in Tables 3 and 4.
Appendix D Qualitative Example
Qualitative analyses for more than 50 popular LMMs are illustrated in Fig. 17-21.
Appendix E Limitations and Future Work
Our MMGenBench-pipeline leverages the powerful capabilities of text-to-image and image representation models. Despite our rigorous efforts to minimize errors, the evaluation pipeline may not achieve complete accuracy for tasks, which presents significant challenges to certain text-to-image and image representation models, such as those involving table images. Our research has the potential to inform and refine future methodologies for modifying target tasks and evaluation techniques, specifically addressing challenges faced by current text-to-image and image representation models. For example, in the context of table images, LaTeX could be utilized to render content and facilitate pixel-level image matching.
In addition, the adoption of the text-to-image model and the image representation model in our MMGenBench-pipeline has led to an increase in evaluation time. Despite this, it is important to note that the MMGenBench-pipeline remains more cost-effective compared to manually constructed benchmarks. Moreover, our MMGenBench-pipeline is fully automated and eliminates the need for human intervention in the evaluation of LMM performance across various domains, making it inherently more scalable than previous benchmarks. This robust scalability enables our pipeline to seamlessly adapt to more powerful models and updated datasets. As the advancements in text-to-image models continue, the time efficiency challenges we currently encounter are expected to improve. For instance, the flux-schnell requires only four steps to generate images, which can significantly accelerate our evaluation process.
Figure 10: A case study of image pattern extraction. GPT-4o is utilized to generate detailed descriptions of images, identify the underlying patterns, and provide explanatory justifications for its identifications.
Model Test Domain
[…]
Qwen2-VL-72B-Instruct [65] 0.553 1.357 0.545 0.710
qwen-vl-max-0809 [6] 0.552 1.363 - -
InternVL2-76B [10] 0.599 1.264 0.599 0.632
GPT4o_HIGH [53] 0.566 1.306 - -
InternVL2-40B [10] 0.566 1.350 0.566 0.696
Ovis1.6-Gemma2-9B [50] 0.582 1.316 0.579 0.667
llava_onevision_qwen2_72b_ov [37] 0.494 1.561 0.491 0.872
llava_onevision_qwen2_72b_si [37] 0.512 1.512 0.514 0.813
Qwen2-VL-7B-Instruct [65] 0.532 1.437 0.524 0.775
InternVL2-26B [10] 0.576 1.320 0.577 0.671
qwen-vl-plus [6] 0.475 1.586 - -
MiniCPM-V-2_6 [75] 0.548 1.386 0.545 0.710
InternVL2-8B [10] 0.547 1.403 0.548 0.701
Ovis1.5-Gemma2-9B [50] 0.521 1.487 0.524 0.808
RBDash_72b [57] 0.525 1.413 0.527 0.740
MMAlaya2 [48] 0.537 1.423 0.541 0.747
Ovis1.5-Llama3-8B [50] 0.526 1.466 0.527 0.795
OmChat [82] 0.439 1.761 0.447 1.028
InternVL-Chat-V1-5 [11] 0.547 1.393 0.547 0.729
llava_onevision_qwen2_7b_si [37] 0.492 1.554 0.497 0.834
XComposer2d5 [11] 0.460 1.646 0.458 0.955
Pixtral-12B [3] 0.552 1.372 0.561 0.686
InternVL2-4B [10] 0.556 1.345 0.556 0.689
llava_onevision_qwen2_7b_ov [37] 0.488 1.590 0.490 0.861
glm-4v-9b [22] 0.511 1.463 0.527 0.730
molmo-7B-D-0924 [14] 0.519 1.490 0.523 0.813
XComposer2_4KHD [11] 0.433 1.781 0.435 1.075
MiniCPM-Llama3-V-2_5 [75] 0.526 1.432 0.530 0.767
VILA1.5-40b [40] 0.483 1.656 0.486 0.958
cambrian_34b [63] 0.546 1.377 0.549 0.701
WeMM [67] 0.488 1.565 0.479 0.888
Llama-3.2-11B-Vision-Instruct [52] 0.526 1.428 0.525 0.768
Qwen2-VL-2B-Instruct [65] 0.487 1.549 0.501 0.806
XComposer2 [10] 0.248 2.890 0.247 2.130
cogvlm2-llama3-chat-19B [26] 0.465 1.523 0.464 0.825
Idefics3-8B-Llama3 [35] 0.460 1.718 0.458 1.004
Mini-InternVL-Chat-4B-V1-5 [21] 0.527 1.433 0.525 0.765
XinYuan-VL-2B-Instruct [71] 0.395 1.951 0.392 1.240
llava_next_yi_34b [43] 0.476 1.636 0.483 0.883
InternVL2-2B [10] 0.476 1.563 0.483 0.848
Phi-3-Vision [1] 0.457 1.678 0.461 0.960
cambrian_13b [63] 0.509 1.526 0.506 0.852
Phi-3.5-Vision [1] 0.479 1.593 0.478 0.892
idefics2_8b [36] 0.449 1.761 0.453 1.015
cambrian_8b [63] 0.512 1.498 0.515 0.810
minimonkey [29] 0.450 1.673 0.455 0.939
Llama-3-MixSenseV1_1 [64] 0.474 1.615 0.477 0.922
llava_next_qwen_32b [43] 0.507 1.501 0.515 0.796
xgen-mm-phi3-interleave-r-v1.5 [73] 0.529 1.454 0.529 0.761
Eagle-X5-13B [59] 0.434 1.812 0.446 1.050
xgen-mm-phi3-dpo-r-v1.5 [73] 0.414 1.671 0.412 0.940
Mini-InternVL-Chat-2B-V1-5 [21] 0.428 1.709 0.440 0.973
llava_next_llama3 [43] 0.476 1.601 0.486 0.881
VILA1.5-13b [40] 0.485 1.695 0.480 1.006
Mantis-8B-siglip-llama3 [31] 0.411 1.833 0.409 1.116
Table 3: Additional results from various LMMs on MMGenBench-Test/Domain. SIM represents SIM-Score, while FID corresponds to FID-Score. The results are arranged in descending order according to the OpenVLM Leaderboard [13].
Model [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Qwen2-VL-72B-Instruct [65] 0.572 0.528 0.502 0.525 0.578 0.559 0.526 0.562 0.563 0.549 0.567 0.603 0.548
qwen-vl-max-0809 [6] 0.568 0.520 0.506 0.536 0.573 0.558 0.521 0.554 0.553 0.541 0.572 0.608 0.538
InternVL2-76B [10] 0.616 0.572 0.540 0.597 0.614 0.611 0.569 0.604 0.594 0.581 0.608 0.640 0.588
GPT4o_HIGH [53] 0.581 0.543 0.516 0.546 0.576 0.574 0.533 0.581 0.567 0.558 0.586 0.620 0.568
InternVL2-40B [10] 0.584 0.527 0.518 0.528 0.592 0.577 0.528 0.581 0.563 0.560 0.574 0.607 0.565
Ovis1.6-Gemma2-9B [50] 0.601 0.547 0.518 0.573 0.593 0.593 0.540 0.595 0.579 0.568 0.580 0.633 0.576
llava_onevision_qwen2_72b_ov [37] 0.512 0.463 0.441 0.463 0.518 0.500 0.458 0.504 0.494 0.478 0.510 0.544 0.509
llava_onevision_qwen2_72b_si [37] 0.530 0.477 0.457 0.500 0.540 0.518 0.467 0.512 0.517 0.515 0.527 0.559 0.509
Qwen2-VL-7B-Instruct [65] 0.550 0.492 0.480 0.514 0.556 0.539 0.488 0.540 0.535 0.515 0.542 0.584 0.520
InternVL2-26B [10] 0.596 0.547 0.522 0.562 0.586 0.583 0.538 0.587 0.572 0.569 0.581 0.627 0.564
qwen-vl-plus [6] 0.497 0.455 0.440 0.386 0.516 0.486 0.457 0.444 0.467 0.464 0.496 0.484 0.486
MiniCPM-V-2_6 [75] 0.559 0.517 0.492 0.524 0.565 0.559 0.500 0.577 0.540 0.547 0.558 0.594 0.522
InternVL2-8B [10] 0.564 0.522 0.501 0.526 0.576 0.557 0.507 0.539 0.545 0.536 0.560 0.596 0.543
Ovis1.5-Gemma2-9B [50] 0.541 0.493 0.470 0.490 0.546 0.526 0.483 0.541 0.535 0.515 0.532 0.557 0.528
RBDash_72b [57] 0.533 0.499 0.491 0.497 0.545 0.531 0.506 0.549 0.540 0.514 0.533 0.553 0.512
MMAlaya2 [48] 0.557 0.511 0.485 0.510 0.554 0.549 0.497 0.554 0.524 0.534 0.549 0.583 0.531
Ovis1.5-Llama3-8B [50] 0.540 0.505 0.485 0.499 0.543 0.526 0.489 0.530 0.539 0.518 0.540 0.565 0.520
OmChat [82] 0.459 0.415 0.394 0.365 0.455 0.437 0.432 0.450 0.457 0.449 0.449 0.468 0.432
InternVL-Chat-V1-5 [11] 0.569 0.522 0.493 0.518 0.562 0.556 0.519 0.556 0.534 0.546 0.559 0.581 0.545
llava_onevision_qwen2_7b_si [37] 0.514 0.462 0.439 0.454 0.520 0.498 0.446 0.494 0.504 0.485 0.514 0.543 0.485
XComposer2d5 [11] 0.446 0.449 0.470 0.411 0.472 0.458 0.472 0.410 0.459 0.441 0.482 0.472 0.478
Pixtral-12B [3] 0.567 0.525 0.513 0.530 0.574 0.557 0.518 0.557 0.554 0.550 0.561 0.579 0.547
InternVL2-4B [10] 0.573 0.524 0.507 0.534 0.578 0.564 0.525 0.540 0.546 0.540 0.574 0.587 0.552
llava_onevision_qwen2_7b_ov [37] 0.504 0.452 0.439 0.470 0.512 0.496 0.432 0.491 0.491 0.465 0.513 0.541 0.478
glm-4v-9b [22] 0.534 0.487 0.474 0.413 0.544 0.522 0.476 0.490 0.490 0.504 0.546 0.536 0.519
molmo-7B-D-0924 [14] 0.537 0.495 0.483 0.447 0.550 0.529 0.496 0.509 0.524 0.507 0.535 0.551 0.518
XComposer2_4KHD [11] 0.444 0.419 0.388 0.387 0.457 0.433 0.425 0.417 0.428 0.435 0.465 0.469 0.437
MiniCPM-Llama3-V-2_5 [75] 0.539 0.498 0.478 0.485 0.547 0.534 0.496 0.523 0.529 0.526 0.539 0.555 0.521
VILA1.5-40b [40] 0.500 0.443 0.417 0.448 0.495 0.492 0.447 0.508 0.458 0.482 0.506 0.536 0.497
cambrian_34b [63] 0.567 0.516 0.503 0.517 0.567 0.552 0.508 0.553 0.535 0.548 0.554 0.592 0.539
WeMM [67] 0.508 0.465 0.445 0.437 0.526 0.493 0.447 0.496 0.495 0.467 0.510 0.519 0.477
Llama-3.2-11B-Vision-Instruct [52] 0.541 0.510 0.501 0.491 0.544 0.529 0.513 0.520 0.513 0.533 0.543 0.561 0.530
Qwen2-VL-2B-Instruct [65] 0.505 0.462 0.439 0.456 0.519 0.489 0.456 0.486 0.495 0.475 0.494 0.539 0.486
XComposer2 [10] 0.264 0.244 0.236 0.175 0.268 0.257 0.207 0.243 0.276 0.255 0.265 0.279 0.249
cogvlm2-llama3-chat-19B [26] 0.486 0.455 0.445 0.382 0.506 0.467 0.440 0.437 0.458 0.461 0.496 0.490 0.464
Idefics3-8B-Llama3 [35] 0.479 0.428 0.403 0.414 0.479 0.465 0.437 0.486 0.447 0.454 0.484 0.511 0.448
Mini-InternVL-Chat-4B-V1-5 [21] 0.541 0.496 0.484 0.488 0.551 0.533 0.489 0.534 0.529 0.522 0.551 0.562 0.525
XinYuan-VL-2B-Instruct [71] 0.393 0.378 0.365 0.337 0.427 0.392 0.387 0.379 0.395 0.393 0.416 0.414 0.398
llava_next_yi_34b [43] 0.498 0.451 0.432 0.391 0.511 0.485 0.429 0.480 0.475 0.472 0.489 0.515 0.461
InternVL2-2B [10] 0.501 0.450 0.440 0.395 0.511 0.479 0.440 0.469 0.479 0.453 0.501 0.498 0.471
Phi-3-Vision [1] 0.469 0.431 0.423 0.365 0.493 0.464 0.442 0.418 0.457 0.449 0.489 0.485 0.473
cambrian_13b [63] 0.535 0.474 0.468 0.465 0.540 0.508 0.477 0.502 0.526 0.521 0.532 0.546 0.519
Phi-3.5-Vision [1] 0.499 0.447 0.428 0.448 0.502 0.485 0.429 0.478 0.474 0.479 0.498 0.524 0.478
idefics2_8b [36] 0.467 0.425 0.408 0.402 0.488 0.451 0.402 0.431 0.443 0.459 0.474 0.486 0.451
cambrian_8b [63] 0.529 0.485 0.460 0.458 0.537 0.520 0.463 0.518 0.518 0.504 0.533 0.541 0.512
minimonkey [29] 0.467 0.422 0.413 0.374 0.488 0.455 0.441 0.430 0.459 0.429 0.478 0.465 0.454
Llama-3-MixSenseV1_1 [64] 0.492 0.448 0.423 0.420 0.489 0.484 0.434 0.476 0.481 0.466 0.490 0.526 0.459
llava_next_qwen_32b [43] 0.526 0.479 0.456 0.449 0.539 0.516 0.468 0.494 0.514 0.489 0.528 0.543 0.501
xgen-mm-phi3-interleave-r-v1.5 [73] 0.549 0.503 0.492 0.495 0.550 0.533 0.515 0.554 0.529 0.511 0.549 0.562 0.524
Eagle-X5-13B [59] 0.446 0.420 0.390 0.387 0.455 0.440 0.407 0.433 0.424 0.440 0.458 0.443 0.439
xgen-mm-phi3-dpo-r-v1.5 [73] 0.420 0.399 0.391 0.289 0.464 0.414 0.402 0.409 0.412 0.430 0.468 0.438 0.415
Mini-InternVL-Chat-2B-V1-5 [21] 0.447 0.408 0.394 0.308 0.474 0.433 0.412 0.410 0.430 0.416 0.455 0.439 0.426
llava_next_llama3 [43] 0.495 0.455 0.427 0.426 0.514 0.485 0.441 0.473 0.478 0.474 0.493 0.516 0.474
VILA1.5-13b [40] 0.499 0.458 0.432 0.473 0.496 0.491 0.468 0.517 0.468 0.484 0.502 0.508 0.486
Mantis-8B-siglip-llama3 [31] 0.424 0.408 0.376 0.365 0.418 0.409 0.411 0.406 0.401 0.433 0.441 0.421 0.422
Table 4: Additional experimental results on MMGenBench-Test, showcasing the SIM-Score for each image pattern. Detailed information about the icons in the first row is available in Fig. 1. The results are sorted in descending order based on the OpenVLM Leaderboard [13].
13 Image Patterns Refer to caption Surreal
This pattern is characterized by its prevalence in depicting scenes that mix elements of fantasy with reality, often creating imaginative or dream-like visuals.
Technology
Highlights themes related to technology, futurism, and modernity, encompassing various aspects of technological advances and speculative futures.
Natural
Encompasses imagery related to natural sceneries or elements, drawing on the beauty and complexity of the natural world.
Artistic
Captures various artistic styles and decorations, reflecting creativity, design elements, and unique art styles.
Color
Focuses on the use and significance of color in images, including aspects like color contrast, schemes, and gradients to convey moods or themes.
Count
Deals with patterns that emphasize numerical aspects, quantities, and the presence of multiple elements, indicating a focus on enumeration or amount.
Orientation
Emphasizes the directionality and alignment within images, capturing how subjects or elements are oriented or directed.
Position
Concerns the relative placement and spatial relationships between elements within an image, highlighting how objects are positioned.
Contextual
Fuses themes of environmental and relational context, shedding light on the settings and the interconnections within images.
Text
Identifies patterns where text is a significant component, showcasing how textual content contributes to the overall message or theme of the image.
Symbol
Centers on the use of symbols and symbolic elements to convey deeper meanings or concepts, drawing from various symbolic traditions.
Geometry
Illustrates a focus on geometric shapes, patterns, and arrangements, underlining the role of structure and form in images.
Motion
Captures the sense of movement or dynamics within images, depicting actions or the concept of motion through visual elements.
Figure 11: $13$ Image Patterns, obtained through GPT-4 Turbo summarization and subsequently verified by humans.
Extracted Image Patterns “Surreal”: 2262, “Structural and Physical Characteristics”: 2016, “Position and Contextual Relationship”: 1775, “Geometry”: 1270, “Orientation and Direction”: 1260, “State and Condition”: 1119, “Text”: 907, “Color”: 420, “Color Contrast”: 404, “Quantity”: 355, “Color Scheme”: 172, “Natural Elements”: 152, “Nature”: 148, “Contextual Relationship”: 137, “Symmetry”: 126, “Texture”: 117, “Anthropomorphism”: 103, “Light and Shadow”: 96, “Fantasy”: 86, “Color Gradient”: 81, “Symbolism”: 80, “Color and Light”: 72, “Lighting”: 70, “Technology”: 67, “Geometric Shapes”: 64, “Contrast”: 64, “Architecture”: 62, “Ornamentation”: 59, “Nature Elements”: 59, “Light and Illumination”: 55, “Lighting and Illumination”: 54, “Color and Texture”: 54, “Reflection”: 49, “Color Pattern”: 49, “Clothing and Accessories”: 46, “Color Palette”: 44, “Color and Contrast”: 40, “Lighting and Color”: 37, “Light and Color”: 37, “Textural Elements”: 35, “Textural”: 34, “Geometric”: 34, “Color Patterns”: 32, “Animal Representation”: 32, “Floral”: 32, “Historical Context”: 31, “Abstract”: 31, “Lighting and Shadow”: 30, “Repetition”: 29, “Fashion”: 29, “Accessories”: 28, “Cultural Elements”: 28, “Geometric Patterns”: 26, “Color and Lighting”: 26, “Character Design”: 25, “Decorative Elements”: 25, “Textural Details”: 25, “Textural and Physical Characteristics”: 24, “Ornamental Design”: 23,
“Medieval and Fantasy Aesthetics”: 1, “Symbolism and Gestures”: 1, “Mechanics and Cybernetics”: 1, “Imagery and Graphics”: 1, “Detail and Complexity”: 1, “Cultural Representations”: 1, “Fusion of Elements”: 1, “Cartoon-like Aesthetic”: 1
Figure 12: Images patterns extracted by GPT-4o and ranked in descending order of frequency.
Prompt for Extraction # Role You are an expert in annotating patterns of images, specializing in the analysis and annotation of images. # Definition Explanation Image Patterns: Refer to repetitive or consistent visual features or structures that appear in images. These patterns can include colors, shapes, textures, arrangements, or even specific objects or symbols. Example Patterns: ”Orientation and Direction”, ”Text”, ”Quantity”, ”Geometry”, ”Surreal”, ”State and Condition”, ”Position and Contextual Relationship”, etc. # Task Instructions 1. Generate a detailed description of the image based on its content, including all details observed in the image. 2. Annotate possible patterns in the image based on the image elements and the description of the image, and explain the reasoning for each identified pattern. # Key Points Carefully inspect all details within the image and annotate possible patterns for the image. Patterns should be based on the contents and visual elements of the image. You may annotate multiple patterns as appropriate. # Output Format A JSON composed of descriptions, patterns, and explanations. The elements are divided into strings, lists of strings, and dictionaries. For example: { ”description”: ”This is a surreal black and white painting depicting a person holding a pile of brains with both hands. The shapes and textures of the brains are clearly visible and exhibit a merging effect. The entire image carries a mysterious and subtly eerie feeling.”, “image_pattern”: [“Text”, “Structural and Physical Characteristics”], “pattern_detail”: { “Text”: “The image contains the text ‘dog’.”, “Structural and Physical Characteristics”: ”The image includes buildings with prominent leaning features.” }}
Figure 13: Prompt for Extraction.
Prompt for Summary # Role You are an expert in image pattern analysis, focusing on the extraction and summarization of image patterns. # Definition Explanation Image Patterns: Refer to repetitive or consistent visual features or structures that appear in images. These patterns can include colors, shapes, textures, arrangements, or even specific objects or symbols. Example Patterns: “Text”, “Quantity”, “Geometry”, “Surreal”, etc. # Task Instructions 1. Summarize the optimal pattern list based on the input JSON data (including image patterns and frequencies). 2. Describe the explanations for summarizing the pattern list. # Key Points Carefully examine the image patterns and their frequencies in the input data, summarize a new list of image patterns, and provide reasons. The new image pattern words should be as short as possible. Not only should you rely on frequency, but also abstract summarization, to abstract from the existing image patterns. # Input Format JSON data, the key is the image pattern, and the value is the frequency the image pattern appears. Example: { “Surreal”: 2262, “Lighting”: 70, “Lighting and Illumination”: 54, } The input data indicates that the image pattern ‘Surreal‘ appears 2262 times, the image pattern ‘Lighting‘ appears 70 times, and the image pattern ‘Lighting and Illumination‘ appears 54 times. # Output Format A JSON composed of patterns and explanations. The elements are divided into lists of strings and dictionaries. Example: { “image_pattern”: [“Surreal”, “Lighting”], “pattern_detail”: { “Surreal”: “Surreal appears frequently”, “Lighting”: “Lighting appears multiple times, and is synonymous with Illumination.” }}
Figure 14: Prompt for Summary.
Prompt for Re-annotation # Role You are an expert in annotating patterns of images, specializing in the analysis and annotation of images. # Definition Explanation Image Patterns: Refer to repetitive or consistent visual features or structures that appear in images. These patterns can include colors, shapes, textures, arrangements, or even specific objects or symbols. Example Patterns: “Surreal”, “Text”, “Count”, etc. # Task Instructions 1. Generate a detailed description of the image based on its content, including all details observed in the image. 2. Annotate possible patterns in the image based on the image elements and the description of the image, and explain the reasoning for each identified pattern. # Key Points 1. Carefully inspect all details within the image and annotate possible patterns for the image. 2. Patterns should be based on the contents and visual elements of the image. You may annotate multiple patterns as appropriate. 3. The labeling pattern can only come from the following patterns: { “Surreal”: “This pattern is characterized by its prevalence in depicting scenes that mix elements of fantasy with reality, often creating imaginative or dream-like visuals.”, “Technology”: “Highlights themes related to technology, futurism, and modernity, encompassing various aspects of technological advances and speculative futures.”, “Natural”: “Encompasses imagery related to natural sceneries or elements, drawing on the beauty and complexity of the natural world.”, “Artistic”: “Captures various artistic styles and decorations, reflecting creativity, design elements, and unique art styles.”, “Color”: “Focuses on the use and significance of color in images, including aspects like color contrast, schemes, and gradients to convey moods or themes.”, “Count”: “Deals with patterns that emphasize numerical aspects, quantities, and the presence of multiple elements, indicating a focus on enumeration or amount.”, “Orientation”: “Emphasizes the directionality and alignment within images, capturing how subjects or elements are oriented or directed.”, “Position”: “Concerns the relative placement and spatial relationships between elements within an image, highlighting how objects are positioned.”, “Contextual”: “Fuses themes of environmental and relational context, shedding light on the settings and the interconnections within images.”, “Text”: “Identifies patterns where text is a significant component, showcasing how textual content contributes to the overall message or theme of the image.”, “Symbol”: “Centers on the use of symbols and symbolic elements to convey deeper meanings or concepts, drawing from various symbolic traditions.”, “Geometry”: “Illustrates a focus on geometric shapes, patterns, and arrangements, underlining the role of structure and form in images.”, “Motion”: “Captures the sense of movement or dynamics within images, depicting actions or the concept of motion through visual elements.” } Where key is the image pattern and value is the explanation of the image pattern. # Output Format A JSON composed of descriptions, patterns, and explanations. The elements are divided into strings, lists of strings, and dictionaries. For example: { “description”: “This is a surreal black and white painting depicting a person holding a pile of brains with both hands. The shapes and textures of the brains are clearly visible and exhibit a merging effect. The entire image carries a mysterious and subtly eerie feeling.”, “image_pattern”: [“Text”, “Count”], “pattern_detail”: { “Text”: “The image contains the text ‘dog’.”, “Count”: “The image contains 3 dogs.” }}
Figure 15: Prompt for Re-annotation.
Evaluation Pipeline Prompt # Role You are an expert in the field of image understanding, focusing on the understanding of images and generating the image caption-prompt. # Definition Explanation image caption-prompt: Refers to the caption or description of an image, used to provide to a Text-to-Image model to generate a new image. Text-to-Image model: Can generate a new image based on the provided image caption-prompt, such as stable diffusion 3, flux, and other image generation models. # Task Description Generate an image caption-prompt based on the input image. # Key Points and Requirements 1. Accurately understand the input image and precisely generate an image caption-prompt. 2. The generated image caption-prompt, when provided to the Text-to-Image model, requires the Text-to-Image model to generate a new image that is as consistent as possible with the input image. 3. The generated image caption-prompt must conform to the preferences of the Text-to-Image model. 4. The generated image caption-prompt should describe the input image in as much detail as possible, and it should be between 20 to 60 words. # Output Format A string, that is the image caption-prompt. No extra output needed.
Figure 16: Evaluation Pipeline Prompt. Text of Multi-modal Input in Figure 3.
Figure 17: More Qualitative Results on MMGenBench. $(1/5)$
Figure 18: More Qualitative Results on MMGenBench. $(2/5)$
Figure 19: More Qualitative Results on MMGenBench. $(3/5)$
Figure 20: More Qualitative Results on MMGenBench. $(4/5)$
Figure 21: More Qualitative Results on MMGenBench. (5/5)55(5/5)( 5 / 5 )